OpenStudy (anonymous):

Well, simplify this: Sqrt(5/4)

OpenStudy (anonymous):

Simplified: square root of 1.25

OpenStudy (lgbasallote):

\[\large \frac{\sqrt 5}{\sqrt 4}\] does that help?

OpenStudy (anonymous):

Yeah, thanks!